If tan(A-B)=1 and sec(A+B)=(2/√3) find the values of A and B and also their general values


 

Grade:12

1 Answers

Tanmay Chaudhari
12 years ago

tan(a-b)=1  means                        sec(a+b)=2/ means

tan(a-b)=tan45 means                   sec(a+b)=sec(30) means

a-b=45 ....(1)                                 a+b=30 ......(2)

solving (1) and (2)

a=37.5 and b=7.5